# Nonlinear moving horizon estimation for robust state and parameter estimation – extended version

### v2

@@ 4 Verifying persistent excitation
In the following, we use the definitions $z:=(x,u,w,p)$ and $\tilde{z}:=(\tilde{x},u,\tilde{w},\tilde{p})$ for any $(x,u,w,p),(\tilde{x},u,\tilde{w},\tilde{p})\in\mathbb{Z}$. In the remainder of this section, we assume that $f$ and $h$ are at least twice continuously differentiable in all of its arguments. Let $z_{s}(s):=z+s(\tilde{z}-z)$ for $s\in[0,1]$ and
@@
[…]
@@
Assumption 12 is motivated by linear systems theory, where detectability is equivalent to the existence of an output injection term which renders the error system asymptotically stable. For any fixed $\eta\in[0,1)$, by using the Schur complement and the definition $\mathcal{Y}(z,\tilde{z}):=PL(z,\tilde{z})$, condition (59) can be transformed into an infinite set of LMIs (linear in the decision variables $P$ and $\mathcal{Y}$). Then, these may be solved under a suitable parameterization of $\mathcal{Y}$ (e.g., polynomial in $z,\tilde{z}$) using a finite set of LMIs and standard convex analysis tools based on semidefinite programming (SDP), e.g., by applying sum-of-squares relaxations [21], by embedding the nonlinear behavior in an LPV model [25], or by suitably gridding the state space and verifying (59) on the grid points (assuming compactness of $\mathbb{Z}$). We also want to emphasize that $L$ does not need to be constant as it is usually required in the context of (adaptive) observer design in order to be able to perform the observer update recursions, cf., e.g., [11]. Instead, the additional degree of freedom resulting from the fact that $L$ may depend on both $z$ and $\tilde{z}$ can be used, e.g., to compensate for nonlinear terms in $A(z,\tilde{z})$ and/or $C(z,\tilde{z})$ from (57). Furthermore, if $L(z,\tilde{z})$ can be chosen such that $\Phi(z,\tilde{z})$ in (58) becomes constant, the condition (59) can be drastically simplified (to one single LMI).
@@
($\left(\{(x_{t},u_{t},w_{t},p)\}_{t=0}^{T-1},\{(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})\}_{t=0}^{T-1}\right)\in\mathbb{Z}^{T}\times\mathbb{Z}^{T}$ (60)
for some $T\in\mathbb{I}_{\geq 0}$, where $x_{t+1}=f(x_{t},u_{t},w_{t},p)$ and $\tilde{x}_{t+1}=f(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})$ for all $t\in\mathbb{I}_{[0,T-1]}$. For the sake of brevity, define
$z_{t}:=(x_{t},u_{t},w_{t},p),\ \tilde{z}_{t}:=(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})$
@@
The following result provides a sufficient condition for the trajectory pair (60) to be an element of the set $\mathbb{E}_{T}$ satisfying Definition 2 based on certain matrix recursions.
@@
$\mathcal{C}_{T}(Z,\tilde{Z}){\,:=}\sum_{t=0}^{T-1}\mu^{T-1-t}\overline{Y}_{t}(z_{t},\tilde{z}_{t})^{\top}\overline{Y}_{t}(z_{t},\tilde{z}_{t}){\succ\,}\alpha I_{o}$ (63)
@@
Proof. Consider the trajectory pair (60) and the outputs $y_{t}=h(x_{t},u_{t},w_{t},p)$ and $\tilde{y}_{t}=h(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})$, $t\in\mathbb{I}_{[0,T-1]}$. Using the mean-value theorem and the definitions from (57), we have that
@@
$\ \ +E(z_{t},\tilde{z}_{t})(p-\tilde{p})$
$\ \ -(\Phi(z_{t},\tilde{z}_{t})Y_{t}+E(z_{t},\tilde{z}_{t})+L(z_{t},\tilde{z}_{t})F(z_{t},\tilde{z}_{t}))(p-\tilde{p}).$
@@
00 0 $=L(z_{t},\tilde{z}_{t})(y_{t}-\tilde{y}_{t}-(y_{t}-\tilde{y}_{t}))$
@@
$M(Z,r):=\max_{\tilde{Z}\in\mathcal{B}(Z,r)}\left\|\frac{\partial\mathcal{C}_{T}}{\partial(Z,\tilde{Z})}(Z,\tilde{Z})\right\|.$
@@
$\|\mathcal{O}_{T}(Z)-\mathcal{C}_{T}(Z,\tilde{Z})\|\geq\|\mathcal{O}_{T}(Z)\|-\|\mathcal{C}_{T}(Z,\tilde{Z})\|.$
@@
$\|\mathcal{C}_{T}(Z,\tilde{Z})\|$ $\geq\|\mathcal{O}_{T}(Z)\|-\|\mathcal{O}_{T}(Z)-\mathcal{C}_{T}(Z,\tilde{Z})\|$
@@
Proposition 15 provides a local result. Applied to the MHE scheme from Section 3, it requires small disturbances $w$ and a good initial guess of ${x}_{0}$ and ${p}_{0}$. This is a standard condition for testing observability properties in the presence of general nonlinear systems, cf., e.g., [30] and [9]. Although the condition on $r$ is not explicitly verifiable in the context of state estimation for general nonlinear systems (due to the fact that $r$ is unknown), checking $\mathcal{O}_{T}(Z)\geq\alpha^{\prime}$ for some $\alpha^{\prime}>0$ yields a reliable heuristic to test in practice if a pair $(Z,\tilde{Z})$ is PE and satisfies (63) or not, which is also evident in the simulation example in Section 5. The construction of $\alpha$ in the proof of Proposition 15 also shows that larger values of $\alpha^{\prime}$ should be chosen if the estimates are more uncertain and therefore $r$ is expected to be large.
@@
Compared to conference version [27], the proposed method to verify PE of a pair of trajectories is a major relaxation in the sense that it can be applied for a much more general class of nonlinear systems; its only limitation lies in its local nature. However, note that global results can be recovered by considering the same class of systems as in [27], i.e., exhibiting the following properties: (i) the dynamics are affine in $p$ and subject to additive disturbances with a linear output map, i.e., $f(w,u,w,p)=\tilde{f}(x,u)+G(x,u)p+Ew$ and $h(x,u,w,p)=Cx+Fw$; (ii) changes in $G(x)$ are directly visible in the output; (iii) the matrix $\Phi$ in (58) is constant by a suitable choice of $L$; (iv) the sets $\mathbb{X},\mathbb{P},\mathbb{W}$ are compact. Then, modifications of the proof of Proposition 14 according to proof of [27, Th. 10] are possible to construct a mapping $\mathcal{C}_{T}$ similar to (63) that satisfies $\mathcal{C}_{T}(Z,\tilde{Z})=\mathcal{C}_{T}(Z,Z)=\mathcal{O}_{T}(Z)$; i.e, such that $\mathcal{C}_{T}(Z,\tilde{Z})$ can be rendered independent of one of its arguments (which would correspond to the unknown true trajectory in the context of state estimation).
@@ 5 Numerical example
which corresponds to the Euler-discretized modified Chua’s circuit system from [33] using the step size $t_{\Delta}=0.01$ under additional disturbances $w\in\mathbb{R}^{4}$ and (noisy) output measurements. The parameters are $b_{1}=12.8$, $b_{2}=19.1$, $a_{1}=0.6$, $a_{2}=-1.1$, $a_{3}=0.45$, which leads to a chaotic behavior of the system. In the following, we treat $w$ as a uniformly distributed random variable with $|w_{i}|\leq 10^{-3},i=1,2,3$ for the process disturbance and $|w_{4}|\leq 0.1$ for the measurement noise. We consider the initial condition $x_{0}=[1,0,-1]^{\top}$ and assume that $x_{t}$ evolves in the (known) set $\mathbb{X}=[-1,3]\times[-1,1]\times[-3,3]$. Furthermore, we consider the case where the exact parameter $a_{3}=:p$ is unknown but contained in the set $\mathbb{P}=[0.2,0.8]$. The objective is to compute the state and parameter estimates $\hat{x}_{t}$ and $\hat{p}_{t}$ by applying the MHE schemes that we proposed in Sections 3.3 and 3 using the initial estimates $\hat{x}_{0}=[-1,0.1,2]^{\top}$ and $\hat{p}_{0}=0.2$.
@@
We construct the i-IOSS Lyapunov function $U$ (Assumption 1) and the set $\mathbb{E}_{N}$ (Definition 2) using the methods from [26, Sec. IV] and Section 4, respectively. The computations are carried out in MATLAB using YALMIP [15] and the semidefinite programming solver MOSEK [18]; details concerning the verification procedure and MHE simulation (including the code) can be found online [29]. We select constant weighting matrices $W_{t}=2P_{\mathrm{p}}$, $V_{t}=100S_{\mathrm{p}}$, $Q_{t}=2(Q_{\mathrm{x}}+Q_{\mathrm{p}})$, and $R_{t}=R_{\mathrm{x}}+R_{\mathrm{p}}$ for all $t\in\mathbb{I}_{\geq 0}$, the discount factors $\eta_{1}=0.934$, and $\eta_{2}=0.9997$, and the horizon length $N=150$. These choices satisfy Assumption 3 and the contraction conditions (24)–(26) (the minimal horizon length for the selected parameters is $N_{\min}=137$). We point out that such a relatively large horizon length is on the one hand due to some conservative steps in the proofs, but on the other hand inherently required here since the dynamics of the system under consideration are rather slow compared to the sampling time $t_{\Delta}$. We consider the proposed MHE scheme presented in Section 3 under two different settings: first, without explicit excitation monitoring and naively assuming uniform PE (satisfaction of Assumption 9), denoted as MHE1; second, with the proposed excitation-dependent regularization update (14) and $\hat{p}_{t}$ in accordance with Remark 7, denoted as MHE2. For the latter, we use the method proposed in Section 4 and check if $X_{t}\in\mathbb{E}_{N_{t}}$ by evaluating $\mathcal{O}_{N_{t}}(Z_{t}^{*})$ defined in (78) at the current optimal solution $Z_{t}^{*}=\big{(}\hat{x}_{t-N_{t}|t}^{*},\hat{p}^{*}_{|t},\{\hat{w}^{*}_{j|t}\}_{j=t-N_{t}}^{t-1}\big{)}$. If the lower bound $\alpha_{t}:=\underline{\lambda}(\mathcal{O}_{T}(Z_{t}^{*}))$ satisfies $\alpha_{t}\geq\alpha$ for the predefined threshold $\alpha=10^{-3}$, we consider that $X_{t}\in\mathbb{E}_{N_{t}}$, and $X_{t}\notin\mathbb{E}_{N_{t}}$ otherwise.
The simulations are carried out in Matlab over $t_{\mathrm{sim}}=5000$ time steps. The estimation errors of MHE1 and MHE2 are shown in Figure LABEL:fig:1, which initially shows a fast convergence of the state and parameter estimation errors for both methods. This is also evident in Figure LABEL:fig:2, which illustrates the respective estimated parameters over time. The initially accurate parameter estimation of both methods results from a sufficiently high excitation level at the beginning of the simulation, which can be seen in Figure LABEL:fig:3.
However, phases of weak excitation occur during the simulation, especially in the time interval $[2000,3000]$ ($\alpha_{t}$ is close to zero in Figure LABEL:fig:3). This renders the parameter unobservable in this interval such that Assumption 9 is violated. Consequently, MHE1 provides very poor parameter estimates in this interval (see the cyan-colored curves in the Figures LABEL:fig:1 and LABEL:fig:2). MHE2, on the other hand, explicitly takes into account the current excitation level and modifies the cost function and parameter estimates accordingly. This efficiently compensates for the phases of poor excitation and leads to significantly better parameter estimation results (see blue curves in Figures LABEL:fig:1 and LABEL:fig:2). In contrast, we can also observe that the unobservability of the parameter has almost no effect on the state estimation results. In fact, the optimal state trajectories produced by MHE1 and MHE2 are almost identical (hence only shown as one red curve in Figure LABEL:fig:1). This can be attributed to the fact that the parameter has a vanishing influence on the cost function when it is unobservable (in particular, the influence is not visible in the fitting error over the horizon). To quantitatively support our statement, we additionally consider the root mean square error (RMSE) for the state and parameter estimation error. As Table 1 shows, the RMSE of the state estimation error are equal for both MHE1 and MHE2, whereas the RMSE of the parameter estimation error for MHE2 is more than six times smaller than MHE1.
Table 1: Comparison of RMSE and $\tau_{\mathrm{avrg}}$ of MHE1 and MHE2.
  Scheme $\mathrm{RMSE}(e_{\mathrm{x}})$ $\mathrm{RMSE}(e_{\mathrm{p}})$ $\tau_{\mathrm{avrg}}\,[\mathrm{ms}]$
  MHE1 0.1707 0.1379 50.01
  MHE2 0.1707 0.0209 49.96
  • $\mathrm{RMSE}(e_{i}):=\sqrt{\frac{1}{N_{\mathrm{sim}}}\sum_{r=0}^{N_{\mathrm{sim}}}\|e_{i,t}\|^{2}}$, $i=\{\mathrm{x},\mathrm{p}\}$
@@
We now compare the average time $\tau_{\mathrm{avrg}}$ required to solve the nonlinear program (4) at each time step (on a standard laptop) for the two schemes MHE1 and MHE2. As the last column in Table 1 shows, these are very similar, which is due to the fact that the additional computation of $\alpha_{t}$ required in MHE2 (which mainly consists of matrix operations) saves time in solving the actual optimization problem due to a better conditioning of the cost function compared to MHE1 (especially due to the regularization update (14)).
@@ 6 Conclusion
We have proposed an MHE scheme to estimate the states and the unknown constant parameters of general nonlinear uncertain discrete time systems. The cost function involves an adaptive regularization term that is adjusted according to the current parameter excitation. We have derived a bound for the state and parameter estimation error that is valid regardless of excitation of the parameter, and in particular also applies if the parameter is never or only rarely excited during operation. Furthermore, we provided a practical method for online excitation monitoring that is applicable to general nonlinear systems. The numerical example has shown that the proposed MHE scheme is able to efficiently compensate for phases of weak excitation and ensures reliable estimation results for all times. An extension of the MHE scheme for time-varying parameters is provided in [28]. Future work might consider component-wise update rules in the MHE designs depending on component-wise excitation monitoring to further improve the estimation performance.

### v3

@@ 4 Verifying persistent excitation
$\textstyle A(z,\tilde{z}):=\int_{0}^{1}\frac{\partial f}{\partial x}(z_{s}(s))ds,\hskip 11.19443pt\textstyle C(z,\tilde{z}):=\int_{0}^{1}\frac{\partial h}{\partial x}(z_{s}(s))ds,$
$\textstyle B(z,\tilde{z}):=\int_{0}^{1}\frac{\partial f}{\partial w}(z_{s}(s))ds,\hskip 8.61108pt\textstyle D(z,\tilde{z}):=\int_{0}^{1}\frac{\partial h}{\partial w}(z_{s}(s))ds,$
$\textstyle E(z,\tilde{z}):=\int_{0}^{1}\frac{\partial f}{\partial p}(z_{s}(s))ds,\hskip 10.33327pt\textstyle F(z,\tilde{z}):=\int_{0}^{1}\frac{\partial h}{\partial p}(z_{s}(s))ds$ (57)
@@
[…]
@@
$\left(\{(x_{t},u_{t},w_{t},p)\}_{t=0}^{T-1},\{(\tilde{x}_{t},u_{t},\tilde{w}_{t},\tilde{p})\}_{t=0}^{T-1}\right)\in\mathbb{Z}^{T}\times\mathbb{Z}^{T}$ (60)
[…]
@@
The following result provides a sufficient condition for the trajectory pair (60) to be an element of the set $\mathbb{E}_{T}$ satisfying Definition 2 based on matrix recursions.
@@
$\mathcal{C}_{T}(Z,\tilde{Z}){\,:=\hskip-1.29167pt}\sum_{t=0}^{T-1}\mu^{T-1-t}\overline{Y}_{t}(z_{t},\tilde{z}_{t})^{\top}\overline{Y}_{t}(z_{t},\tilde{z}_{t}){\succ\,}\alpha I_{o}$ (63)
@@
[…]
@@
[…]
@@
$0$ $=L(z_{t},\tilde{z}_{t})(y_{t}-\tilde{y}_{t}-(y_{t}-\tilde{y}_{t}))$
@@
[…]
@@
[…]
@@
[…]
@@
Proposition 15 provides a local result. Applied to the MHE scheme from Section 3, it requires small disturbances $w$ and a good initial guess of ${x}_{0}$ and ${p}_{0}$. This is a standard condition for testing observability properties in the presence of general nonlinear systems, cf., e.g., [31] and [9]. Although the condition on $r$ is not explicitly verifiable in the context of state estimation for general nonlinear systems (due to the fact that $r$ is unknown), checking $\mathcal{O}_{T}(Z)\geq\alpha^{\prime}$ for some $\alpha^{\prime}>0$ yields a reliable heuristic to test in practice if a pair (Z,Z~)($\tilde{Z}$) is PE and satisfies (63) or not, which is also evident in the simulation example in Section 5. The construction of $\alpha$ in the proof of Proposition 15 also shows that larger values of $\alpha^{\prime}$ should be chosen if the estimates are more uncertain and therefore $r$ is expected to be large.
@@
[…]
@@ 5 Numerical example
which corresponds to the Euler-discretized modified Chua’s circuit system from [34] using the sampling interval $t_{\Delta}=0.01$ under additional disturbances $w\in\mathbb{R}^{4}$ and (noisy) output measurements. The parameters are $b_{1}=12.8$, $b_{2}=19.1$, $a_{1}=0.6$, $a_{2}=-1.1$, $a_{3}=0.45$, which leads to a chaotic behavior of the system. In the following, we treat $w$ as a uniformly distributed random variable with $|w_{i}|\leq 10^{-3},i=1,2,3$ for the process disturbance and $|w_{4}|\leq 0.1$ for the measurement noise. We consider the initial condition $x_{0}=[1,0,-1]^{\top}$ and assume that $x_{t}$ evolves in the (known) set $\mathbb{X}=[-1,3]\times[-1,1]\times[-3,3]$. Furthermore, we consider the case where the exact parameter $a_{3}=:p$ is unknown but contained in the set $\mathbb{P}=[0.2,0.8]$. The objective is to compute the state and parameter estimates $\hat{x}_{t}$ and $\hat{p}_{t}$ by applying the MHE scheme proposed in Section 3 using the initial estimates $\hat{x}_{0}=[-1,0.1,2]^{\top}$ and $\hat{p}_{0}=0.2$.
@@
We construct the i-IOSS Lyapunov function $U$ (Assumption 1) and the set $\mathbb{E}_{N}$ (Definition 2) using the methods from [27, Sec. IV] and Section 4, respectively. The computations are carried out in MATLAB using YALMIP [15] and the semidefinite programming solver MOSEK [18]; details concerning the verification procedure and MHE simulation (including the code) can be found online [30]. We select constant weighting matrices $W_{t}=2P_{\mathrm{p}}$, $V_{t}=100S_{\mathrm{p}}$, $Q_{t}=2(Q_{\mathrm{x}}+Q_{\mathrm{p}})$, and $R_{t}=R_{\mathrm{x}}+R_{\mathrm{p}}$ for all $t\in\mathbb{I}_{\geq 0}$, the discount factors $\eta_{1}=0.934$, and $\eta_{2}=0.9997$, and the horizon length $N=150$. These choices satisfy Assumption 3 and the contraction conditions (24)–(26) (the minimal horizon length for the selected parameters is $N_{\min}=137$). While the relatively high value of the horizon length indicates conservatism in our results, it also seems natural here, as the sampling interval $t_{\Delta}$ is relatively small compared to the dynamics of the system. In the following, we consider the MHE scheme presented in Section 3 with two different settings: first, without explicit excitation monitoring by naively assuming uniform PE (Assumption 9); second, using the proposed excitation-dependent adaptive regularization and selecting the parameter estimate $\hat{p}_{t}$ in accordance with Remark 7. For the latter, we use the method proposed in Section 4 and check if $X_{t}\in\mathbb{E}_{N_{t}}$ by evaluating $\mathcal{O}_{N_{t}}(Z_{t}^{*})$ defined in (78) at the current optimal solution $Z_{t}^{*}=\big(\hat{x}_{t-N_{t}|t}^{*},\hat{p}^{*}_{|t},\{\hat{w}^{*}_{j|t}\}_{j=t-N_{t}}^{t-1}\big)$. If the lower bound $\alpha_{t}:=\underline{\lambda}(\mathcal{O}_{T}(Z_{t}^{*}))$ satisfies $\alpha_{t}\geq\alpha$ for the predefined threshold $\alpha=10^{-3}$, we consider that $X_{t}\in\mathbb{E}_{N_{t}}$, and $X_{t}\notin\mathbb{E}_{N_{t}}$ otherwise.
Figure 1: Estimation results of the proposed MHE scheme with adaptive regularization (red) compared to naive MHE without excitation motoring (blue) and two MHE designs for state estimation only, relying on inaccurate (fixed) parameters $\hat{p}=\hat{p}_{0}=0.2$ (yellow) and $\hat{p}=0.5$ (purple). In the top plot, the red and blue curves coincide.
The estimation results are shown Figure 1. Here, we first note that MHE can lead to significant state estimation errors under parametric uncertainties if incorrect (fixed) parameters are used (compare the yellow and purple curves), which highlights the necessity and effectiveness of estimating states and parameters jointly in a combined MHE framework. In particular, both the naive and adaptive (joint) MHE schemes show rapid convergence of the state and parameter estimation errors, see the blue and red curves. Here, the high accuracy of the parameter estimates at the beginning (see the second and third plot) results from a sufficiently high excitation (compare the bottom plot). During the simulation, however, there are also phases of weak excitation—especially in the time interval $[2000,3000]$, where the excitation level $\alpha_{t}$ is close to zero. This essentially renders the parameter unobservable, thereby violating Assumption 9. As a result, the naive MHE scheme provides poor parameter estimates in this interval, which can be clearly seen from the blue curve in the second and third plots in Figure 1. In contrast, MHE with the proposed adaptive regularization explicitly takes the excitation level into account and is thus able to efficiently compensate for phases of weak excitation, resulting in significantly better parameter estimation results (see the red curves).
While parameter unobservability must obviously be addressed appropriately to ensure reliable parameter estimates, it has almost no impact on the state estimates (the blue and red curves coincide in the top plot in Figure 1). In fact, the state trajectories produced by the naive and the adaptive MHE schemes are almost identical, which can be attributed to the fact that the parameter has a vanishing influence on the cost function when it is unobservable. This is also evident in Table 1, where we provide the root mean square error (RMSE) for the state and parameter estimates. Specifically, the RMSE of the state estimation errors is equal for both MHE schemes (their difference lies in the order of magnitude of $10^{-15}$), whereas the RMSE of the parameter estimation error for MHE using adaptive regularization is more than six times smaller than the naive MHE scheme without excitation monitoring.
Table 1: Naive and adaptive MHE for state and parameter estimation.
MHE setup $\mathrm{RMSE}(e_{\mathrm{x}})$ $\mathrm{RMSE}(e_{\mathrm{p}})$ $\tau_{\mathrm{avrg}}\,[\mathrm{ms}]$
naive 0.1707 0.1379 50.01
adaptive 0.1707 0.0209 49.96
  • $\mathrm{RMSE}(e_{i}){\,:=\,}\sqrt{\frac{1}{N_{\mathrm{sim}}}\sum_{t=0}^{N_{\mathrm{sim}}}\|e_{i,t}\|^{2}}$, $i=\{\mathrm{x},\mathrm{p}\}$, $N_{\mathrm{sim}}{\,=\,}5000$
@@
We now compare the average time $\tau_{\mathrm{avrg}}$ required to solve the nonlinear program in (4) at each time step $t$ for MHE with and without adaptive regularization using a standard laptop. Interestingly, the last column in Table 1 shows that the computation times are very similar. This can be attributed to the fact that the additional computation of $\alpha_{t}$ required for the adaptation mechanism (which mainly consists of matrix operations) saves time in solving the actual optimization problem due to a better conditioning of the cost function compared to naive MHE, especially due to the regularization update (14).
@@ 6 Conclusion
In this paper, we have proposed an MHE scheme to estimate the states and the unknown constant parameters of general uncertain nonlinear discrete-time systems. The cost function involves an adaptive regularization term that is adjusted according to the current parameter excitation. We have derived a bound for the state and parameter estimation error that is valid regardless of excitation of the parameter, and in particular also applies if the parameter is never or only rarely excited during operation. Furthermore, we provided a practical method for online excitation monitoring that is applicable to general nonlinear systems. The numerical example has shown that the proposed MHE scheme is able to efficiently compensate for phases of weak excitation and ensures reliable estimation results for all times. An extension of the MHE scheme for time-varying parameters is provided in [29]. Future work might consider component-wise adaptation mechanisms in MHE depending on component-wise excitation monitoring to further improve the overall estimation performance.